# Trustworthy Transfer Learning: A Survey

### v1

@@ 4.4.2 Uncertainty Quantification
Uncertainty quantification is essential for decision-making and optimization in machine learning and artificial intelligence (?). For example, high-stakes applications such as medical diagnostics (?) and autonomous driving (?) require both accurate class predictions and quantification of prediction uncertainty. Generally, there are two types of prediction uncertainty (?): aleatoric (data) uncertainty involving the inherent randomness and variability in the data, and epistemic (model) uncertainty caused by a lack of knowledge about the optimal model parameters. These uncertainties can be formally explained using the Bayesian posterior distribution (?).
Definition 10 (Aleatoric and Epistemic Uncertainty (?))
@@
[…]
@@
The calibration of uncertainty estimates is vital in determining the trustworthiness of model outputs. A well-calibrated model should provide accurate predictions when it is confident and indicate high uncertainty when it is likely to be incorrect. Thus, calibration can be considered as an orthogonal metric for accuracy when evaluating machine learning systems. In particular, ? (?) conduct a systematic evaluation of traditional uncertainty quantification models under distribution shifts. They observe that the quality of uncertainty consistently degrades with increasing distribution shifts between source and target domains. To solve this problem, various frameworks have been proposed to improve uncertainty quantification under distribution shifts across domains.
@@
    Temperature Scaling: ? (?) derive an upper bound on the expected calibration error in the target domain in terms of the importance-weighted classification error and the error of a domain discriminator. Building on the idea of temperature scaling (?), they propose a calibration algorithm by minimizing the upper bound over source and target samples. Similarly, ? (?) develop an adaptive importance weighting approach with lower bias and variance of the estimated calibration errors to improve the uncertainty quantification under the covariate shift assumption. Instead, ? (?) focus on learning two calibration functions based on a real in-distribution calibration set and a synthetic out-of-distribution calibration set respectively, and then adaptively combine the two calibrators. ? (?) optimize the calibration objective function (i.e., temperature scaling optimization (?)) using a labeled pseudo-target set created via mixup (?) over pseudo-labeled target samples.
@@
    Conformal Prediction: The model predicts a set of labels instead of a single label (?, ?, ?). Assuming the true importance weights (e.g., $w^{*}(x,y)=p_{T}(x,y)/p_{S}(x,y)$) are known, ? (?) and ? (?) study the weighted conformal predictions under covariate shifts and label shifts, respectively. Based on jackknife+ (?), ? (?, ?) formulate the sampling weighted jackknife+ prediction interval to handle covariate shifts with finite-sample coverage guarantee. ? (?) design prediction sets that are robust against all distribution shifts with bounded $f$-divergence. ? (?, ?) further investigate prediction sets in an online setting where the data distribution can shift continuously over time. In addition, ? (?) construct probably approximately correct (PAC) prediction sets under bounded covariate shifts in the scenarios with known importance weights and an uncertainty set of possible importance weights. The follow-up work (?) constructs prediction sets with PAC guarantees in the presence of label shifts. The crucial idea is to compute confidence intervals of importance weights (?) through Gaussian elimination.
@@
    Bayesian Learning: ? (?) develop an approximate Bayesian inference approach based on posterior regularization that captures the distribution difference between source and target domains. ? (?) study the uncertainty quantification problem in the context of test-time adaptation. They develop a probabilistic graphical model for covariate shift scenarios, followed by an instantiated ensemble approach to estimate the uncertainty of trained models over test samples. In addition, transferable Gaussian processes (?, ?, ?, ?, ?) can be applied to model uncertainty in the target domain by leveraging knowledge from the source domain.
4.5 Other Trustworthiness Concerns
4.5.1 Accountability and Auditability
Accountability is crucial to evaluate the trustworthiness of AI outcomes, as it identifies the organizations and individuals responsible for these results. More specifically, ? (?) defines accountability as “a relationship between an actor and a forum, in which the actor has an obligation to explain and to justify his or her conduct, the forum can pose questions and pass judgment, and the actor may face consequences”. ? (?) further analyzes five key aspects of this definition, including the actor, the forum, the relationship between them, the content and criteria of the account, and the potential consequences resulting from the account. In this case, auditability refers to systematic evaluations to guarantee accountability (?). Given the remarkable performance of fine-tuned large language models (LLMs), auditing LLMs has been studied through different principled assessments (?, ?). Recently, ? (?) discuss data and AI model markets that facilitate the sharing, discovery, and integration of data and AI models among multiple parties. These markets can enhance knowledge transfer between pre-trained AI models and user-specific tasks, but they raise fundamental concerns regarding accountability in these systems. Further research can be conducted to guarantee accountability for transfer learning systems in supporting model and data knowledge sharing.
4.5.2 Sustainability and Environmental Well-being
To establish the trustworthiness of machine learning and artificial intelligence systems, it is crucial to evaluate resource usage and energy consumption within their entire supply chain (?, ?). Notably, ? (?) introduce a simple notion of computational cost in producing AI results.
Definition 11 (Cost of an AI Result (?))
@@ 4.5.2 Sustainability and Environmental Well-being
$\mathrm{Cost}(R)\propto E\cdot D\cdot H$ (24)
@@
To reduce the computational cost, green AI (?, ?, ?) has been promoted by improving the efficiency of AI models with positive impacts on the environment. Several efficiency metrics have been introduced, including carbon emission, electricity usage, floating-point operations (FLOPs), elapsed runtime, and the number of parameters. Transfer learning techniques demonstrated significant improvements in training efficiency by leveraging knowledge from pre-trained models (?, ?). This is because these approaches reduce (1) the size of training data $D$ and hyperparameters $H$, and (2) the number of trainable model parameters (via parameter-efficient fine-tuning (?, ?)). Furthermore, ? (?) recently propose a GreenTrainer method to minimize the FLOPs of LLM fine-tuning via adaptive backpropagation. ? (?) take a first look into the carbon footprint of federated learning, by quantifying carbon emissions from hardware training and communication between server and clients. In real-world applications, transfer learning has been applied to lower energy consumption and reduce carbon emissions by reusing pre-trained models (?). For example, ? (?) and ? (?) analyze transfer learning techniques for garbage classification and flower classification in terms of both prediction accuracy and carbon emissions.
5 Applications
@@
Trustworthy transfer learning has been widely applied to artificial intelligence and machine learning fields, including computer vision (?), natural language processing (?), and graph learning (?). In addition, this section highlights real-world applications of trustworthy transfer learning in scientific discovery.
5.1 Agriculture
@@
Transfer learning techniques have been applied to various precision agriculture applications (?, ?). Specifically, to improve the management of agricultural stakeholders, ? (?) and ? (?) propose process-guided machine learning frameworks, which transfer knowledge from simulated data generated by soil-vegetation radiative transfer modeling to real-world field data for precise monitoring of cover crop traits. ? (?) analyze the transferability of support vector regression models for estimating leaf nitrogen concentration across different plant species. Besides, pre-trained vision models have been fine-tuned for crop mapping (?), crop pest classification (?), and plant phenotyping (?).
5.2 Bioinformatics
@@
Notably, ? (?) have introduced an attention-based foundation model Geneformer pre-trained on over 30 million single-cell transcriptomes to capture network dynamics (e.g., gene interactions). They also demonstrate the effectiveness of Geneformer in various downstream tasks with limited data through fine-tuning. Later, ? (?) illustrate the efficacy of the pre-trained large language model GPT-4 in cell type annotation of single-cell RNA-seq data. Besides, ? (?) and ? (?) develop unified transfer learning frameworks for open-world single-cell annotation across different species and tissues, as well as for batch correction and multi-omics integration. Similarly, ? (?) study the clustering of single-cell RNA-seq data on the small disease- or tissue-specific data sets by leveraging prior knowledge from large reference data sets. ? (?) and ? (?) leverage architecture surgery based transfer learning techniques to understand cellular heterogeneity.
@@
In addition, recent efforts (?, ?, ?) have been devoted to protein representation learning for downstream tasks using language models pre-trained on a large protein corpus. Typically, ? (?) introduce a protein transfer learning benchmark TAPE for learning transferable protein representation, while ? (?) further improve the quality of protein representation by considering the geometry of representation space. ? (?) also exploit the pre-trained ProteinMPNN model (?) to extract embeddings of input proteins, which are then used to predict stability changes for protein point mutations.
5.3 Healthcare
@@
Transfer learning advances the development of effective and efficient health care services (?). For example, ? (?) develop a federated transfer learning framework for privacy-preserving wearable healthcare systems (e.g., Parkinson’s disease auxiliary diagnosis). ? (?) and ? (?) further understand the impact of the source domain/model on the downstream medical imaging tasks in the context of transfer learning. To enforce health equity across ethnic groups, ? (?), ? (?), and ? (?) propose transferring knowledge from majority groups with sufficient data to minority groups with limited data. In addition, transfer learning techniques have been applied to drug discovery (?). Specifically, ? (?) propose a functional rationalized meta-learning algorithm to enable knowledge transfer across assays for virtual screening and ADMET prediction. ? (?) and ? (?) adopt pre-training and fine-tuning strategies for molecular property prediction and drug-target interaction prediction, respectively.
5.4 Education
@@
Transfer learning has been studied in Educational Data Mining (EDM) for predicting student performance in higher education (?). Over the past decade, Massive Open Online Courses (MOOCs) have supported millions of learners around the world. Early predictions of student performance are crucial for enabling timely interventions in these courses. Transfer learning has been explored to predict student performance in ongoing courses by leveraging knowledge from previous courses. To be specific, ? (?) leverage knowledge from both previous courses and previous weeks of the same course to make real-time predictions for learners in MOOCs. Instead of relying on handcrafted features, ? (?) aim to learn domain-invariant representation by using an auto-encoder and correlation alignment (?) between source and target courses. Similarly, ? (?) study the transferability of early success prediction models across MOOCs from different domains and topics. Besides, ? (?) explore the transferability of student performance in addressing the cold-start problem for new courses in intelligent tutoring systems. More recently, large language models such as GPT-4 and ChatGPT have gained significant attention for improving instructional efficiency and student engagement (e.g., by creating interactive homework with feedback and follow-up questions) (?, ?).
5.5 Robotics
@@
Sim-to-real transfer aims to transfer knowledge from simulation to real-world environments when training reinforcement learning models for robotic learning (?, ?). Recently, this framework has been studied in various robotic learning tasks, including Rubik’s cube (?), human pose estimation (?), vision-and-language navigation (?), biped locomotion (?), etc. Specifically, various strategies have been proposed to address the domain shift between simulation and real-world environments (?, ?, ?, ?). One is to use distribution alignment regularization to learn domain-invariant representation (?, ?). Another strategy is domain randomization (?, ?), which aims to train the model using a diverse set of randomized simulated environments, rather than relying on a single simulated environment. ? (?) and ? (?) further theoretically highlight the benefits of domain randomization for sim-to-real transfer.
5.6 E-commerce
@@
Cross-domain recommendation aims to generate reliable recommendations in a target domain by exploiting knowledge from source recommender systems. It has been studied from various perspectives, e.g., matrix factorization (?, ?), neural collaborative filtering (?, ?, ?), graph neural network (?, ?), large language models (?), etc. In addition to prediction accuracy, recent works also investigate the trustworthiness properties of cross-domain recommender systems, such as adversarial vulnerability (?) and user privacy (?).
6 Open Questions and Future Trends
@@
6.1 Benchmarking Negative Transfer
Negative transfer can roughly be defined as the phenomenon (?) where “transferring knowledge from the source can have a negative impact on the target learner”.
Definition 12 (Negative Transfer (?, ?))

### v2

@@ 4.4.2. Uncertainty Quantification
Uncertainty quantification is essential for decision-making and optimization in machine learning and artificial intelligence (naeini2015obtaining). For example, high-stakes applications such as medical diagnostics (begoli2019need) and autonomous driving (michelmore2020uncertainty) require both accurate class predictions and quantification of prediction uncertainty. Generally, there are two types of prediction uncertainty (hullermeier2021aleatoric): aleatoric (data) uncertainty involving the inherent randomness and variability in the data, and epistemic (model) uncertainty caused by a lack of knowledge about the optimal model parameters. These uncertainties can be formally explained using the Bayesian posterior distribution (ChanAQS20).
Definition 4.4 (Aleatoric and Epistemic Uncertainty (ChanAQS20)).
@@
(23) $\underbrace{p(y^{*}|x^{*},D)}_{\text{Total uncertainty}}=\int\underbrace{p(y^{*}|x^{*},f)}_{\text{Aleatoric uncertainty}}\cdot\underbrace{p(f|D)}_{\text{Epistemic uncertainty}}df$
@@
The calibration of uncertainty estimates is vital in determining the trustworthiness of model outputs. A well-calibrated model should provide accurate predictions when it is confident and indicate high uncertainty when it is likely to be incorrect. Thus, calibration can be considered as an orthogonal metric for accuracy when evaluating machine learning systems. In particular, (SnoekOFLNSDRN19) conduct a systematic evaluation of traditional uncertainty quantification models under distribution shifts. They observe that the quality of uncertainty consistently degrades with increasing distribution shifts between source and target domains. To solve this problem, various frameworks have been proposed to improve uncertainty quantification under distribution shifts across domains.
@@
    Temperature Scaling: (ParkBWL20) derive an upper bound on the expected calibration error in the target domain in terms of the importance-weighted classification error and the error of a domain discriminator. Building on the idea of temperature scaling (guo2017calibration), they propose a calibration algorithm by minimizing the upper bound over source and target samples. Similarly, (TransCal) develop an adaptive importance weighting approach with lower bias and variance of the estimated calibration errors to improve the uncertainty quantification under the covariate shift assumption. Instead, (zou2023adaptive) focus on learning two calibration functions based on a real in-distribution calibration set and a synthetic out-of-distribution calibration set respectively, and then adaptively combines the two calibrators. (hu2024pseudocalibration) optimize the calibration objective function (i.e., temperature scaling optimization (guo2017calibration)) using a labeled pseudo-target set created via mixup (zhang2018mixup) over pseudo-labeled target samples.
@@
    Conformal Prediction: The model predicts a set of labels instead of a single label (romano2020classification; lei2018distribution; angelopoulos2021gentle). Assuming the true importance weights (e.g., $w^{*}(x,y)=p_{T}(x,y)/p_{S}(x,y)$) are known, (tibshirani2019conformal; PodkopaevR21) study the weighted conformal predictions under covariate shifts and label shifts, respectively. Based on jackknife+ (Barber2019PredictiveIW), (prinster2022jaws; prinster2023jaws) formulate the sampling weighted jackknife+ prediction interval to handle covariate shifts with a finite-sample coverage guarantee. (cauchois2024robust) design prediction sets that are robust against all distribution shifts with bounded $f$-divergence. (gibbs2021adaptive; gibbs2024conformal) further investigate prediction sets in an online setting where the data distribution can shift continuously over time. In addition, (park2022pac) construct probably approximately correct (PAC) prediction sets under bounded covariate shifts in scenarios with known importance weights and an uncertainty set of possible importance weights. The follow-up work (si2024pac) constructs prediction sets with PAC guarantees in the presence of label shifts. The crucial idea is to compute confidence intervals of importance weights (lipton2018detecting) through Gaussian elimination.
@@
    Bayesian Learning: (ChanAQS20) develop an approximate Bayesian inference approach based on posterior regularization that captures the distribution difference between source and target domains. (zhou2021bayesian) study the uncertainty quantification problem in the context of test-time adaptation. They develop a probabilistic graphical model for covariate shift scenarios, followed by an instantiated ensemble approach to estimate the uncertainty of trained models over test samples. In addition, transferable Gaussian processes (yu2007gaussian; bonilla2007multi; cao2010adaptive; maddox2021fast; DINO) can be applied to model uncertainty in the target domain by leveraging knowledge from the source domain.
4.5. Other Trustworthiness Concerns
4.5.1. Accountability and Auditability
Accountability is crucial to evaluate the trustworthiness of AI outcomes, as it identifies the organizations and individuals responsible for these results. More specifically, (bovens2007analysing) defines accountability as “a relationship between an actor and a forum, in which the actor has an obligation to explain and to justify his or her conduct, the forum can pose questions and pass judgment, and the actor may face consequences”. (wieringa2020account) further analyze five key aspects of this definition, including the actor, the forum, the relationship between them, the content and criteria of the account, and the potential consequences resulting from the account. In this case, auditability refers to systematic evaluations to guarantee accountability (raji2020closing). Given the remarkable performance of fine-tuned large language models (LLMs), auditing LLMs has been studied through different principled assessments (mokander2023auditing; amirizaniani2024auditllm). Recently, (pei2023data) discuss data and AI model markets that facilitate the sharing, discovery, and integration of data and AI models among multiple parties. These markets can enhance knowledge transfer between pre-trained AI models and user-specific tasks, but they raise fundamental concerns regarding accountability in these systems. Further research can be conducted to guarantee accountability for transfer learning systems in supporting model and data knowledge sharing.
4.5.2. Sustainability and Environmental Well-being
To establish the trustworthiness of machine learning and artificial intelligence systems, it is crucial to evaluate resource usage and energy consumption within their entire supply chain (nikolinakos2023ethical; budennyy2022eco2ai). Notably, (schwartz2020green) introduce a simple notion of computational cost in producing AI results.
Definition 4.5 (Cost of an AI Result (schwartz2020green)).
@@ 4.5.2. Sustainability and Environmental Well-being
(24) $\mathrm{Cost}(R)\propto E\cdot D\cdot H$
@@
To reduce the computational cost, green AI (schwartz2020green; huang2024towards; memmel2024position) has been promoted by improving the efficiency of AI models with positive impacts on the environment. Several efficiency metrics have been introduced, including carbon emission, electricity usage, floating-point operations (FLOPs), elapsed runtime, and the number of parameters. Transfer learning techniques demonstrated significant improvements in training efficiency by leveraging knowledge from pre-trained models (YosinskiCBL14; he2019rethinking). This is because these approaches reduce (1) the size of training data $D$ and hyperparameters $H$, and (2) the number of trainable model parameters (via parameter-efficient fine-tuning (Adapters; LoRA)). Furthermore, (huang2024towards) recently propose a GreenTrainer method to minimize the FLOPs of LLM fine-tuning via adaptive backpropagation. (qiu2023first) take a first look into the carbon footprint of federated learning, by quantifying carbon emissions from hardware training and communication between server and clients. In real-world applications, transfer learning has been applied to lower energy consumption and reduce carbon emissions by reusing pre-trained models. For example, (kunwar2024managing; ahmed2024flowers) analyze transfer learning techniques for garbage classification and flower classification in terms of both prediction accuracy and carbon emissions.
5. Applications
@@
Trustworthy transfer learning has been widely applied to artificial intelligence and machine learning fields, including computer vision (NeyshaburSZ20), natural language processing (ding2023parameter), and graph learning (WNN). In addition, this section highlights real-world applications of trustworthy transfer learning in scientific discovery.
5.1. Agriculture
@@
Transfer learning techniques have been applied to various precision agriculture applications (ma2024transfer; adve2024aifarms). Specifically, to improve the management of agricultural stakeholders, (zhang2021transfer; wang2023airborne) propose process-guided machine learning frameworks, which transfer knowledge from simulated data generated by soil-vegetation radiative transfer modeling to real-world field data for precise monitoring of cover crop traits. (wan2022combining) analyze the transferability of support vector regression models for estimating leaf nitrogen concentration across different plant species. Besides, pre-trained vision models have been fine-tuned for crop mapping (jo2022towards), crop pest classification (thenmozhi2019crop), and plant phenotyping (sama2023new).
5.2. Bioinformatics
@@
Notably, (Theodoris2023TransferLE) introduce an attention-based foundation model Geneformer pre-trained on over 30 million single-cell transcriptomes to capture network dynamics (e.g., gene interactions). They also demonstrate the effectiveness of Geneformer in various downstream tasks with limited data through fine-tuning. Later, (hou2024assessing) illustrate the efficacy of the pre-trained large language model GPT-4 in cell type annotation of single-cell RNA-seq data. Besides, (hu2020iterative) develop a unified transfer learning framework for open-world single-cell classification across different species and tissues. Similarly, (mieth2019using) study the clustering of single-cell RNA-seq data on the small disease- or tissue-specific data sets by leveraging prior knowledge from large reference data sets. (hetzel2022predicting; lotfollahi2022mapping) leverage architecture surgery based transfer learning techniques to understand cellular heterogeneity.
@@
In addition, recent efforts (rao2019evaluating; detlefsen2022learning; heinzinger2019modeling) have been devoted to protein representation learning for downstream tasks using language models pre-trained on a large protein corpus. Typically, (rao2019evaluating) introduce a protein transfer learning benchmark TAPE for learning transferable protein representation, and (detlefsen2022learning) further improve the quality of protein representation by considering the geometry of representation space. (dieckhaus2024transfer) also exploit the pre-trained ProteinMPNN model (dauparas2022robust) to extract embeddings of input proteins, which are then used to predict stability changes for protein point mutations.
5.3. Healthcare
@@
Transfer learning advances the development of effective and efficient health care services (jayaraman2020healthcare). For example, (FedHealth) develop a federated transfer learning framework for privacy-preserving wearable healthcare systems (e.g., Parkinson’s disease auxiliary diagnosis). (raghu2019transfusion; matsoukas2022makes) further understand the impact of the source domain/model on the downstream medical imaging tasks in the context of transfer learning. To enforce health equity across ethnic groups, (gao2020deep; toseef2022reducing; lee2023reducing) propose transferring knowledge from majority groups with sufficient data to minority groups with limited data. In addition, transfer learning techniques have been applied to drug discovery (cai2020transfer). Specifically, (yao2021functionally) propose a functional rationalized meta-learning algorithm to enable knowledge transfer across assays for virtual screening and ADMET prediction. (goh2018using; dalkiran2023transfer) adopt pre-training and fine-tuning strategies for molecular property prediction and drug-target interaction prediction, respectively.
5.4. Education
@@
Transfer learning has been studied in Educational Data Mining (EDM) for predicting student performance in higher education (hunt2017transfer). Over the past decade, Massive Open Online Courses (MOOCs) have supported millions of learners around the world. Early predictions of student performance are crucial for enabling timely interventions in these courses. Transfer learning has been explored to predict student performance in ongoing courses by leveraging knowledge from previous courses. To be specific, (boyer2015transfer) leverage knowledge from both previous courses and previous weeks of the same course to make real-time predictions for learners in MOOCs. Instead of relying on handcrafted features, (ding2019transfer) aim to learn domain-invariant representation by using an auto-encoder and correlation alignment (CORAL) between source and target courses. Similarly, (SwamyMK22) study the transferability of early success prediction models across MOOCs from different domains and topics. Besides, (Schmucker2022transferable) explore the transferability of student performance in addressing the cold-start problem for new courses in intelligent tutoring systems. More recently, large language models such as GPT-4 and ChatGPT have gained significant attention for improving instructional efficiency and student engagement (e.g., by creating interactive homework with feedback and follow-up questions) (vanzo2024gpt; Kasneci2023ChatGPTFG).
5.5. Robotics
@@
Sim-to-real transfer aims to transfer knowledge from simulation to real-world environments when training reinforcement learning models for robotic learning (dai2024acdc; peng2018sim). Recently, this framework has been studied in various robotic learning tasks, including Rubik’s cube (akkaya2019solving), human pose estimation (DoerschZ19), vision-and-language navigation (AndersonSTMPBL20), biped locomotion (YuKTL19), etc. Specifically, various strategies have been proposed to address the domain shift between simulation and real-world environments (tzeng2020adapting; peng2018sim; pinto2017robust; rusu2017sim). One is to use distribution alignment regularization to learn domain-invariant representation (tzeng2020adapting; DIRL). Another strategy is domain randomization (andrychowicz2020learning; tobin2017domain), which aims to train the model using a diverse set of randomized simulated environments, rather than relying on a single simulated environment. (chen2022understanding; hu2023provable) further theoretically highlight the benefits of domain randomization for sim-to-real transfer.
5.6. E-commerce
@@
Cross-domain recommendation aims to generate reliable recommendations in a target domain by exploiting knowledge from source recommender systems. It has been studied from various perspectives, e.g., matrix factorization (man2017cross; samra2024cross), neural collaborative filtering (HuZY18; kanagawa2019cross; li2020ddtcdr), graph neural network (GRADE; zhao2019cross), large language models (petruzzelli2024instructing), etc. In addition to prediction accuracy, recent works also investigate the trustworthiness properties of cross-domain recommender systems, such as adversarial vulnerability (chen2019data) and user privacy (yang2024federated).
6. Open Questions and Future Trends
@@
6.1. Benchmarking Negative Transfer
Negative transfer can roughly be defined as the phenomenon (TL_survey_2010) where “transferring knowledge from the source can have a negative impact on the target learner”.
Definition 6.1 (Negative Transfer (TL_survey_2010; WangDPC19)).